Convert into a mole.
 22 g of Carbon dioxide.
Convert into a mole.
 22 g of Carbon dioxide.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given mass of Carbon dioxide = 22 g
Molar mass of carbon dioxide (CO2) = (1 × 12) + (2 × 16)
= 12 + 32 = .44 g
Mole of carbon dioxide =22 = 0.5 mole
44

Thanks